OpenStudy (korosh23):

Absolute value function has the form of this: f(x)=|ax+b| The x-intercept is (3/2,0) and y-intercept is (0,6) Find the value of a and b.

OpenStudy (korosh23):

\[{x \epsilon R}\]

OpenStudy (korosh23):

That is the domain and our range is {yI y> or equal to 0, y is also real number} Find the value of a and b

OpenStudy (korosh23):

can you help?

OpenStudy (korosh23):

We know this is in the form of Y=mx + b

OpenStudy (korosh23):

m and a are the same. These two indicate the slope. b is the y-intercept. I know how to find b, but not a

OpenStudy (anonymous):

It's not actually of the form mx+b.

OpenStudy (korosh23):

ok

OpenStudy (anonymous):

Let's consider an example. Help me graph this function: f(x) = |2x +6|

OpenStudy (anonymous):

What is y when x =0? What is x when y =0?

OpenStudy (korosh23):

Ok first we find the x and y intercepts

OpenStudy (korosh23):

y= (0,6) x= (-3,0)

OpenStudy (korosh23):

Then we graph

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

How about y when x = -6?

OpenStudy (korosh23):

x= (-6,0)?

OpenStudy (anonymous):

No, I mean if you plug in x = -6, what is f(x)? f(-6) = |2(-6)+6|

OpenStudy (korosh23):

oh I am sorry y= 6 absolute value of a negative number is postive number

OpenStudy (jdoe0001):

\(\large { f(x)={\color{blue}{ y}}=|a{\color{brown}{ x}}+b|\qquad \begin{cases} (\frac{3}{2}&,0)\\ {\color{brown}{ x}}&{\color{blue}{ y}} \\\hline\\ (0&,6)\\ {\color{brown}{ x}}&{\color{blue}{ y}} \end{cases}\implies \begin{cases} {\color{blue}{ 0}}=\left|a\cdot {\color{brown}{ \frac{3}{2}}}+b\right| \\ \\\hline\\ {\color{blue}{ 6}}=|a\cdot {\color{brown}{ 0}}+b| \end{cases} \\ \quad \\ 6=|a\cdot 0+b|\implies 6=|b| \\ \quad \\ \textit{keep in mind that }|a+b|\iff |a|+|b| }\)

OpenStudy (anonymous):

Good! So we get those three points and we can graph it from that. (-6,6), (-3,0), (0, 6) 

OpenStudy (korosh23):

ok @jdoe0001 then how do you find a using that way?

OpenStudy (anonymous):

Now my question is this: What did the +3 in our equation do? And what did the 2 in our equation do?

OpenStudy (korosh23):

@SmoothMath I do not understand what do you mean exactly by +3 and 2 ?

OpenStudy (anonymous):

Our function was f(x) = |2x+3| what did the +3 do to the graph?

OpenStudy (korosh23):

3 is the y-intercept? 2 is the slope?

OpenStudy (anonymous):

Close. But look at the y-axis. Does it cross at 3? No, it crosses at 6. The y-intercept is 6.

OpenStudy (jdoe0001):

@korosh23 well... does it matter how you're meant to get it? graphicaly, or usnig a system of equations? if graphically, then SmoothMath way would work

OpenStudy (anonymous):

You're correct about the slope. 2 made the graph have a steepness of 2.

OpenStudy (jdoe0001):

@korosh23 anyhow, in my case, I'm using a system of equations, but either method will do, just a matter of what is it the exercise is covering

OpenStudy (korosh23):

No I want it algebrically. Well I know the answer, just can I explain to you guys and see if it is correct or not?

OpenStudy (jdoe0001):

well, post what you have, or take a screenshot of it on paper, and post it away :)

OpenStudy (korosh23):

ok wait a moment

OpenStudy (korosh23):

Absolute value function has the form of this: f(x)=|ax+b| The x-intercept is (3/2,0) and y-intercept is (0,6) b is 6 since it is the y-intercept. a is the slope and I found it this way. 

OpenStudy (korosh23):

@jdoe0001 @SmoothMath

OpenStudy (anonymous):

You should have stuck with me on my explanation =)

OpenStudy (anonymous):

b is not the y-intercept, as I showed you in my example.

OpenStudy (anonymous):

Go look at it again. b is not the y- intercept. What is it?

OpenStudy (korosh23):

but math can be applied differently, In my textbook it says b is the y-intercept and also my teacher admit that too.

OpenStudy (jdoe0001):

I'm afraid I'm biased towards the algebraic approach but the graphic one will do @korosh23 :)

OpenStudy (korosh23):

@jdoe0001 that means I am correct?

OpenStudy (anonymous):

Haha it's not a question of bias. It's a question of fact.

OpenStudy (jdoe0001):

hmmm not sure what you got for "a" or "b" @korosh23 so.... can't say is wrong or right